OpenStudy (anonymous):

Helen can paint a living room in 5 hours. If Bill help Helen they can paint the room in 3 hours. How long would it take Bill to paint the living room alone?

OpenStudy (anonymous):

I know that you set it up like Helen=1/5 Bill=1/t Total=1/3 So i think it's like 1/5+1/t=1/3 but i don't know if you find a common denominator next or what to do after that/: please help

OpenStudy (anonymous):

that is a good start

OpenStudy (anonymous):

\[\frac{1}{5}+\frac{1}{t}=\frac{1}{3}\] \[\frac{1}{t}=\frac{1}{3}-\frac{1}{5}=\frac{2}{15}\] \[t=\frac{15}{2}=7.5\]

OpenStudy (anonymous):

Thank you so much !

OpenStudy (anonymous):

yw